# Mamba-ND: Selective State Space Modeling for Multi-Dimensional Data

### v3

@@ 1 Introduction
Figure 1: Mamba-ND  outperforms Transformers while significantly reducing the number of parameters. On ImageNet-1k, we compare against ViT [11]. On HMDB-51[23], we compare against Video-Swin [30]. On ERA5, we compare against Cli-ViT [38].
@@ 5.2 Video Action Recognition
We sample 32 frames at a frame interval of 2 from each video clip, which amounts to around 2 seconds of video. We use the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $45e-$ 4. The learning rate on the backbone is multiplied by a factor of 0.1. We train our model with a global batch size of 64 for 50 epochs. We select Inception-I3D [4] and ConvNeXt-I3D [31] as our convolutional baseline. For transformer-based models, we select Video Swin Transformer [30]. We report the results in Tab. 2. Our method achieves a $+2.8$ accuracy compared to Video Swin Transformer while using only 39% of the parameters.
@@ 5.3 Global Weather Forecasting
Table 3: ERA5 5.625°normal-$\degree$ Weather Forecasting Results. Anomaly Correlation Coefficient (ACC) and RMSE on geopotential at the 500 hPa level. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
@@
Mamba-3D-UNETR-S Mamba 36M 81.5 8G [16, 16, 16]
@@
Mamba-3D-UNETR-B Mamba 107M 84.1 9G [16, 16, 16]
@@ 8 Acknowledgement
8.1 Erratas
We correct the following differences of model performance in the main paper: on HMDB-51, our method achieves a +2.8 (instead of +0.9) accuracy compared to Video Swin Transformer and +7.4 (instead of +5.5) accuracy compared to ConvNeXT. We also correct the learning rate to be 5.0e-4 (instead of 1.0e-3) for the main results. For completeness, we provide results under different learnig rate in Tab. 7. While learning rate affects model performance by a non-trival margin, our model consistently outperforms competitors regardless of learning rates.
Table 7: Ablation Studies on Learning Rate
Ablation Study on Learning Rate. We report top-1 accuracy on the HMDB-51 dataset of swpeeing learning rate between 1.0e-3 and 1.0e-4. Learning rate HMDB-51$\uparrow$ 1.0e-3 59.0 5.0e-4 60.9 1.0e-4 60.2

### v5

@@ 1 Introduction
Figure 1: Mamba-ND  outperforms Transformers while significantly reducing the number of parameters. On ImageNet-1k, we compare against ViT [13]. On HMDB-51[29], we compare against Video-Swin [37]. On ERA5, we compare against Cli-ViT. On BTCV, we compare against UNETR[19]. [45].
@@ 5.1 Image Classification
Table 3: Video Classification Results on Kinetics400 and Breakfast Classification Results. For Kinectics400, we report results using 32 frames (/32) and 64 frames (/64).
[…]
Swin-T 4×3 78.8 28M
Swin-S 4×3 80.6 49M
Swin-B 4×3 80.6 88M
ViViT-L 4×3 80.6 310M
TimeSformer-L 1×3 80.7 121M
Mamba-3D/32 4×3 80.9 38M
Mamba-3D/64 4×3 81.9 39M
[…]
Distant Supervision TimeSformer 89.9
ViS4mer Swin+SSM 88.2
TranS4mer SSM 90.3
Mamba-3D SSM 91.2
@@ 5.2 Video Action Recognition
We sample 32 frames at a frame interval of 2 from each video clip, which amounts to around 2 seconds of video. We use the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $46e-$ 4. The learning rate on the backbone is multiplied by a factor of 0.1. We train our model with a global batch size of 64 for 50 epochs. We select Inception-I3D [6] and ConvNeXt-I3D [38] as our convolutional baseline. For transformer-based models, we select Video Swin Transformer [37],TimeSfromerr [4], and ViViT[2]. We conducted experiments of UCF-101[50], HMDB-51[29] and Kinects-400[26] datasets. All models are initialized with the ImageNet-1K pretrained checkpoints. We report the results in Tab. 2 and Tab. 3. Our method achieves a $+1.3$ accuracy on Kinectics-400 dataset compared to Video Swin Transformer while using only 44% of the parameters.
To further explore the capabilities of Mamba-3D on video understanding task, we evaluate Mamba-3D on long video classification task on Breakfast[28] dataset. Mamba-3D outperforms previous state-of-the art models based on Transformers (Distant Supervision[34]), SSMs (TranS4mer[25]), and hybrid architectures (ViS4mer[24]).
@@ 5.3 Global Weather Forecasting
Table 4: ERA5 5.625°$\degree$ Weather Forecasting Results. Anomaly Correlation Coefficient (ACC) and RMSE on geopotential at the 500 hPa level. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
@@
Mamba-3D-UNETR-S Mamba 36M 83.1 8G [16, 16, 16]
@@
Mamba-3D-UNETR-B Mamba 107M 84.7 9G [16, 16, 16]
@@ 9.7 Ordering
10 Implementation Details
10.1 Model Size
Mamba-2D-S has 24 layers and a hidden size of 384. Mamba-2D-B has 24 layers and a hidden size of 768. Mamba-3D-S has 32 layers and a hidden size of 384. Mamba-3D-S+ has 32 layers and a hidden size of 512. Mamba-3D-B has 32 layers and a hidden size of 768. For ERA5 weather forecast, we use a 12 layer version of Mamba-3D-B. In terms of the parameter count, two Mamba Layers is roughly equivalent to one ViT block of the same dimension.
10.2 Additional Hyperparameters
On HMDB-51, we use a learning rate of 6.0e-4. On UCF-101, we use a learning rate of 1.0e-3. We set RandAug to $(9,4)$ for both experiments. We do not use label smoothing or cutmix. We set the dropout and drop path rate both to 0.1. On ImageNet-1K, we use a learning rate of 1.0e-3. We set RandAug to (2, 10). We set label smoothing to 0.1, Mixup to 0.8 and CutMix to 1.0. We set the dropout and drop path rate both to 0.1. For all image and video ablations, we fix the learning rate to be 1e-3 and use a patch size of 16. On BTCV, we use a learning rate of 1.0e-4 and a drop rate of 0.1. On ERA5, we use a learning rate of 5.0e-4 and a drop rate of 0.1.
11 Limitations
While we have provided a relatively comprehensive set of experiments to explore the design space of multi-dimensional selective state-space models, there remain unexplored areas due to the exponential number of possible orderings. In this paper, we primarily explored various combinations of row-major ordering. However, there may be other reasonable choices such as diagonal or zig-zag patterns. We leave these possibilities for future work.